OpenStudy (anonymous):

A certain keyboard has 5 more than twice as many keys across as it does vertically. The special keyboard has a total of 168 keys. ((((((Find the amount of keys positioned vertically and across)))))

OpenStudy (anonymous):

168=(2a+5)??????????

OpenStudy (anonymous):

168=.5(a)(2a+5) ???????

jimthompson5910 (jim_thompson5910):

let x = # of keys across and y = # of keys vertically

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

"A certain keyboard has 5 more than twice as many keys across as it does vertically" that all means x = 2y + 5

jimthompson5910 (jim_thompson5910):

The special keyboard has a total of 168 keys tells us that x*y = 168

OpenStudy (akashdeepdeb):

Multiply vertical with across = 168! Gte quadratic equation solve it, then get the answer!

jimthompson5910 (jim_thompson5910):

x*y = 168 (2y + 5)*y = 168 ... plug in x = 2y + 5 2y^2 + 5y = 168 2y^2 + 5y - 168 = 0 I'll let you finish

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

so we are = 0 right now

jimthompson5910 (jim_thompson5910):

you now have to solve 2y^2 + 5y - 168 = 0 for y

jimthompson5910 (jim_thompson5910):

use the quadratic formula

OpenStudy (anonymous):

8??

jimthompson5910 (jim_thompson5910):

good y = 8, now use this to find x

OpenStudy (anonymous):

did you get across = 21 vertical = 8 ????????

jimthompson5910 (jim_thompson5910):

yes I did

OpenStudy (anonymous):

great

jimthompson5910 (jim_thompson5910):

notice how 21*8 = 168

OpenStudy (anonymous):

bam